# AI-Powered Energy Algorithmic Trading: Integrating Hidden Markov Models with Neural Networks

### v6

@@ Abstract
In quantitative finance, machine-learning methods are key for alpha generation. This study introduces a refreshing approach that combines Hidden Markov Models (HMM) and neural networks integrated with Black-Litterman portfolio optimization. The approach was tested during the COVID period (2019-2022), where it achieved an 83% return with a Sharpe ratio of 0.77. Two risk models were incorporated to enhance risk management, particularly during the volatile periods. The methodology was implemented on the QuantConnect platform, which was chosen for its robust framework and experimental reproducibility. The system is designed to predict future price movements and includes a three-year warm-up period to ensure the proper use of the algorithm. It focuses on highly liquid, large-cap energy stocks to ensure stable and reliable results, while also accounting for broker payments. The dual-model alpha system utilizes log returns to select the optimal state based on historical performance. It combines state predictions with neural network outputs derived from historical data to generate trading signals. This study presents an in-depth examination of the trading system’s architecture, data pre-processing, training, and performance. The full code and backtesting data are available under QuantConnect’s terms.
keywords:
Hidden Markov Models; Neural Networks; Algorithmic Trading; Machine Learning; Energy Sector; Portfolio OptimizationJEL Codes: C45, C53, G11, G17
@@ 1 Introduction
Algorithmic trading influences financial markets by enabling the rapid and exact execution of trading strategies beyond human capability Dananjayan et al. (2023). Machine learning integration has further reinvented this field by providing novel techniques for pattern recognition and predictive modeling Afua et al. (2024).
@@
This study outlines a novel methodology that fuses the HMM and neural networks to create a dual-model alpha generation system. The aim is to leverage the HMM’s ability to capture temporal dependencies and market regimes, along with the power of neural networks to learn subtle patterns from historical price data Oelschlager and Adam (2020); El Morr et al. (2022). The strategy uses Black-Litterman portfolio optimization combined with two risk management models.
@@
The main objective is to develop an adaptable trading strategy that predicts price fluctuations and optimizes trading decisions. The results indicate the potential of this unified approach, with the algorithm achieving a 83% return and a Sharpe ratio of 0.77 during the COVID period (2019-2022).
@@ 2 Background
HMMs, deep learning, and multi-model AI are crucial for optimizing algorithmic trading strategies. HMMs simulate systems with partially observable states, whereas deep learning utilizes neural networks to capture complex patterns. Multi-model AI integrates various machine learning models to improve robustness and accuracy Giudici et al. (2024).
The use of HMMs in financial markets has evolved from simple regime-switching models to advanced volatility modeling methods Gorynin et al. (2017). Deep learning has progressed from basic neural networks to sophisticated architectures, such as LSTM, CNN, and transformer-based models Shiri et al. (2023). Multi-model AI, employing ensemble methods, surpasses single ML algorithms in terms of robustness Aelgani and Vadlakonda (2023). These theories form the basis for advanced trading strategies. HMMs provide insights into market regimes and volatility, deep learning models forecast price movements, and multi-model AI enhances the overall performance.
@@ 2.1 Previous Research
Previous research has highlighted the utility of HMMs in identifying market regimes and in volatility modeling. In addition, deep learning techniques have shown potential for time-series forecasting and sentiment analysis Lara-Benitez et al. (2021). Multi-model AI approaches, such as stacking and boosting, have improved prediction accuracy across various fields Odegua (2019). HMM studies often use regime-switching models and probabilistic analysis, whereas deep learning focuses on neural network architectures and training methods. Multi-model AI research combines different algorithms through ensemble techniques to leverage their unique strengths.
@@ 2.2 Challenges
Despite their advantages, HMMs face challenges in capturing the market complexity. Simple HMMs may not ideally interpret market dynamics because of their inability to capture both immediate and extended trends Oelschlager and Adam (2021). Deep learning models capture nonlinear relationships but require substantial data and computational resources. Although they often outperform traditional methods, their benefits can be modest and context-dependent, especially when data are limited or computational costs are high Jiang (2020). Multi-model AI approaches offer enhanced robustness and predictive performance but are more complex to implement Wang et al. (2021). Variability in data, methods, and implementation can lead to inconsistencies, complicating the reproduction of results Chen et al. (2022).
@@ 2.3 Research Gaps and Future Directions
Researchers often downplay the benefits of HMMs, deep learning, and multi-model AI in trading strategies owing to the lack of standardized frameworks for strategy standardization and data pre-processing. This absence challenges the creation of single and multi-model systems Afua et al. (2024b). Bridging these gaps is essential to craft trading strategies capable of traversing complex financial markets. Future studies should explore integrating techniques to capture both linear and nonlinear market behavior, and investigate the impact of real-time market conditions on model performance Bharath et al. (2023).
@@ 2.4 Tools and Hypotheses
QuantConnect facilitates advanced trading by supporting multi-model deployment using the LEAN open-source engine. It standardizes data and algorithm development, improving efficiency and reproducibility, while addressing current limitations Rashid (2010).
@@ 3 Methodology
This section explains how the dual-model system for generating trading signals was created using the HMM and neural networks. The QuantConnect platform was used in this project because it provides significant market data and streamlines strategy testing and deployment. Since 2013, it has been refined through over two million algorithms, ensuring reproducibility and transparency QuantConnect (2024).
@@ 3.1 Universe Selection
This study used a dynamic universe with two filters to select tradable securities. The first filter evaluates liquidity through the dollar volume by selecting stocks with the highest total traded value. The second filter focuses on energy sector stocks based on market capitalization and identifies the top 20 companies in this sector. Larger companies typically exhibit more stable and predictable performance, beneficial for analysis Sauberschwarz and Weiss (2017). This way, this dynamic universe prioritizes liquidity and market capitalization. Establishing a solid foundation for scientific models to operate effectively without artificial complexity Rakovic (2018).
@@ 3.3 Portfolio Construction
Proper portfolio construction is crucial for the success of the algorithm. The Black-Litterman model was chosen for its innovative approach, combining expected returns and risk. Created by Fischer Black and Robert Litterman, this model merges investor views with market equilibrium for better asset allocation Cayirli (2019).
@@
This model increases algorithm modularity, separating alpha generation from portfolio management Tzang et al. (2020). Unlike signal-based portfolio construction, it reduces the probability of inaccurate signals, preventing inefficient management in achieving the optimal Sharpe ratio. The Black-Litterman model improves system reliability by clearly dividing alpha generation and portfolio management. It directs alpha model insights into risk management and execution models and minimizes unnecessary trades and risks.
@@ 3.4 Risk Management and Execution Models
3.4.1 Maximum Drawdown Model
@@
3.4.2 Trailing Stop Model
@@ 4.4 Advantages of the Approach
Utilizing the HMM alongside neural networks offers notable advantages for predictive modeling in financial markets. HMMs capture temporal dependencies crucial for sequential data like stock prices, while neural networks excel in recognizing complex patterns and relationships Alwateer et al. (2023); Qamar and Zardari (2023). By combining these approaches, a more complete and accurate prediction model is achieved, leveraging the strengths of both strategies. This hybrid approach enhances the robustness by reducing the biases inherent in any single model, resulting in more reliable trading signals. Risk management strategies also play a crucial role in guarding against potential losses and improving the overall portfolio stability. The adaptability gained from merging HMMs and neural networks is essential for maintaining a high performance across different market conditions.
@@ 4.5 Future Research Directions
4.5.1 Wavelet Transform for Enhanced Market Pattern Recognition
One potential approach is to apply a wavelet transform to the pre-processing strategy of the data. Wavelet transforms capture both time and frequency domain information, improving the potential of the model to recognize patterns in price data, which is crucial to ensuring its success Kwon et al. (2022). This is vital as it accounts for the dynamic nature of the market, making it more effective than static methods such as Pearson’s correlation coefficient.
4.5.2 Monte Carlo Simulations for Model Generalizability
The use of extensive Monte Carlo simulations across diverse temporal frames is another promising area of research. These simulations help analyze the generalization capabilities of the model, evaluating its reliability and identifying possible cases of overfitting or incorrect market predictions Zillich et al. (2009). This ensures the stability and accuracy of the algorithmic trading system in a real trading scenario.
4.5.3 Hyperparameter Tuning and Ensemble Methods
Employing novel neural network architectures along with automated hyperparameter fine-tuning can enhance forecasting capabilities by finding the optimal architecture for a specific dataset Wu et al. (2024). Additionally, incorporating an ensemble machine learning method for HMM and neural network systems can enable the model to adapt dynamically to market fluctuations, improving overall alpha generation performance Ong and Ahmad (2011).
4.5.4 Diversifying Equities for Growth and Stability
@@ 5 Conclusion
The combined use of the HMM and neural networks in a dual-model approach has shown success in enhancing the accuracy and resilience of trading signals. During the COVID period (2019-2022), the dual-model system achieved an 83% return and a Sharpe ratio of 0.77, highlighting its practical potential.
@@
Use of AI tools declaration
The author declares that artificial intelligence tools were employed only to assist in drafting initial versions of certain sections of this manuscript. All AI generated content has been meticulously revised and rewritten by the author. The author takes absolute responsibility for the final content of this article.

### v7

@@ Abstract
In quantitative finance, machine-learning methods are key for alpha generation. This study introduces a refreshing approach that combines Hidden Markov Models (HMM) and neural networks integrated with Black-Litterman portfolio optimization. The approach was tested during the COVID period (2019–2022), where it achieved an 83% return with a Sharpe ratio of 0.77. Two risk models were incorporated to enhance risk management, particularly during the volatile periods. The methodology was implemented on the QuantConnect platform, which was chosen for its robust framework and experimental reproducibility. The system is designed to predict future price movements and includes a three-year warm-up period to ensure the proper use of the algorithm. It focuses on highly liquid, large-cap energy stocks to ensure stable and reliable results, while also accounting for broker payments. The dual-model alpha system utilizes log returns to select the optimal state based on historical performance. It combines state predictions with neural network outputs derived from historical data to generate trading signals. This study presents an in-depth examination of the trading system’s architecture, data pre-processing, training, and performance. The full code and backtesting data are available under QuantConnect’s terms.
Keywords: Hidden Markov Models; Neural Networks; Algorithmic Trading; Machine Learning; Energy Sector; Portfolio Optimization
JEL Codes: C45, C53, G11, G17
@@ 1 Introduction
Algorithmic trading influences financial markets by enabling the rapid and exact execution of trading strategies beyond human capability [dananjayan2023]. Machine learning integration has further reinvented this field by providing novel techniques for pattern recognition and predictive modeling [afua2024].
@@
This study outlines a novel methodology that fuses the HMM and neural networks to create a dual-model alpha generation system. The aim is to leverage the HMM’s ability to capture temporal dependencies and market regimes, along with the power of neural networks to learn subtle patterns from historical price data [oelschlager2020, elmorr2022]. The strategy uses Black-Litterman portfolio optimization combined with two risk management models.
@@
The main objective is to develop an adaptable trading strategy that predicts price fluctuations and optimizes trading decisions. The results indicate the potential of this unified approach, with the algorithm achieving a 83% return and a Sharpe ratio of 0.77 during the COVID period (2019–2022).
@@ 2 Background
HMMs, deep learning, and multi-model AI are crucial for optimizing algorithmic trading strategies. HMMs simulate systems with partially observable states, whereas deep learning utilizes neural networks to capture complex patterns. Multi-model AI integrates various machine learning models to improve robustness and accuracy [giudici2024].
The use of HMMs in financial markets has evolved from simple regime-switching models to advanced volatility modeling methods [gorynin2017]. Deep learning has progressed from basic neural networks to sophisticated architectures, such as LSTM, CNN, and transformer-based models [shiri2023]. Multi-model AI, employing ensemble methods, surpasses single ML algorithms in terms of robustness [aelgani2023]. These theories form the basis for advanced trading strategies. HMMs provide insights into market regimes and volatility, deep learning models forecast price movements, and multi-model AI enhances the overall performance.
@@ 2.1 Previous Research
Previous research has highlighted the utility of HMMs in identifying market regimes and in volatility modeling. In addition, deep learning techniques have shown potential for time-series forecasting and sentiment analysis [larabenitez2021]. Multi-model AI approaches, such as stacking and boosting, have improved prediction accuracy across various fields [odegua2019]. HMM studies often use regime-switching models and probabilistic analysis, whereas deep learning focuses on neural network architectures and training methods. Multi-model AI research combines different algorithms through ensemble techniques to leverage their unique strengths.
@@ 2.2 Challenges
Despite their advantages, HMMs face challenges in capturing the market complexity. Simple HMMs may not ideally interpret market dynamics because of their inability to capture both immediate and extended trends [oelschlager2021]. Deep learning models capture nonlinear relationships but require substantial data and computational resources. Although they often outperform traditional methods, their benefits can be modest and context-dependent, especially when data are limited or computational costs are high [jiang2020]. Multi-model AI approaches offer enhanced robustness and predictive performance but are more complex to implement [wang2021]. Variability in data, methods, and implementation can lead to inconsistencies, complicating the reproduction of results [chen2022].
@@ 2.3 Research Gaps and Future Directions
Researchers often downplay the benefits of HMMs, deep learning, and multi-model AI in trading strategies owing to the lack of standardized frameworks for strategy standardization and data pre-processing. This absence challenges the creation of single and multi-model systems [afua2024b]. Bridging these gaps is essential to craft trading strategies capable of traversing complex financial markets. Future studies should explore integrating techniques to capture both linear and nonlinear market behavior, and investigate the impact of real-time market conditions on model performance [bharath2023].
@@ 2.4 Tools and Hypotheses
QuantConnect facilitates advanced trading by supporting multi-model deployment using the LEAN open-source engine. It standardizes data and algorithm development, improving efficiency and reproducibility, while addressing current limitations [rashid2010].
@@ 3 Methodology
Note: This figure illustrates the complete workflow from data collection through signal generation to portfolio execution.
This section explains how the dual-model system for generating trading signals was created using the HMM and neural networks. The QuantConnect platform was used in this project because it provides significant market data and streamlines strategy testing and deployment. Since 2013, it has been refined through over two million algorithms, ensuring reproducibility and transparency [quantconnect2024].
@@ 3.1 Universe Selection
This study used a dynamic universe with two filters to select tradable securities. The first filter evaluates liquidity through the dollar volume by selecting stocks with the highest total traded value. The second filter focuses on energy sector stocks based on market capitalization and identifies the top 20 companies in this sector. Larger companies typically exhibit more stable and predictable performance, beneficial for analysis [sauberschwarz2017]. This way, this dynamic universe prioritizes liquidity and market capitalization. Establishing a solid foundation for scientific models to operate effectively without artificial complexity [rakovic2018].
@@ 3.2.1 Model Architecture
Note: The model uses five hidden states to capture different market regimes including bull, bear, and transitional phases.
@@
Note: The feedforward network consists of four hidden layers with ReLU activations, trained using the Adam optimizer.
@@ 3.3 Portfolio Construction
Proper portfolio construction is crucial for the success of the algorithm. The Black-Litterman model was chosen for its innovative approach, combining expected returns and risk. Created by Fischer Black and Robert Litterman, this model merges investor views with market equilibrium for better asset allocation [cayirli2019].
@@
This model increases algorithm modularity, separating alpha generation from portfolio management [tzang2020]. Unlike signal-based portfolio construction, it reduces the probability of inaccurate signals, preventing inefficient management in achieving the optimal Sharpe ratio. The Black-Litterman model improves system reliability by clearly dividing alpha generation and portfolio management. It directs alpha model insights into risk management and execution models and minimizes unnecessary trades and risks.
@@ 3.4 Risk Management and Execution Models
Maximum Drawdown Model
@@
Trailing Stop Model
@@ 4.1 Key Statistics
Note: Performance metrics cover the full backtesting period from January 2019 to January 2022, encompassing the COVID-19 market volatility.
@@ 4.2 Detailed Trading Metrics
Note: The low trading frequency (40 total orders) reflects the conservative, signal-based approach of the dual-model system.
@@ 4.3 Cumulative Returns
Note: The light blue line represents the backtested strategy while the gray line shows the S&P 500 benchmark. Notable outperformance occurred during the COVID-19 market volatility in May 2020 and September 2021.
@@ 4.4 Advantages of the Approach
Utilizing the HMM alongside neural networks offers notable advantages for predictive modeling in financial markets. HMMs capture temporal dependencies crucial for sequential data like stock prices, while neural networks excel in recognizing complex patterns and relationships [alwateer2023, qamar2023]. By combining these approaches, a more complete and accurate prediction model is achieved, leveraging the strengths of both strategies. This hybrid approach enhances the robustness by reducing the biases inherent in any single model, resulting in more reliable trading signals. Risk management strategies also play a crucial role in guarding against potential losses and improving the overall portfolio stability. The adaptability gained from merging HMMs and neural networks is essential for maintaining a high performance across different market conditions.
@@ 4.5 Future Research Directions
Wavelet Transform for Enhanced Market Pattern Recognition
One potential approach is to apply a wavelet transform to the pre-processing strategy of the data. Wavelet transforms capture both time and frequency domain information, improving the potential of the model to recognize patterns in price data, which is crucial to ensuring its success [kwon2022]. This is vital as it accounts for the dynamic nature of the market, making it more effective than static methods such as Pearson’s correlation coefficient.
Monte Carlo Simulations for Model Generalizability
The use of extensive Monte Carlo simulations across diverse temporal frames is another promising area of research. These simulations help analyze the generalization capabilities of the model, evaluating its reliability and identifying possible cases of overfitting or incorrect market predictions [zillich2009]. This ensures the stability and accuracy of the algorithmic trading system in a real trading scenario.
Hyperparameter Tuning and Ensemble Methods
Employing novel neural network architectures along with automated hyperparameter fine-tuning can enhance forecasting capabilities by finding the optimal architecture for a specific dataset [wu2024]. Additionally, incorporating an ensemble machine learning method for HMM and neural network systems can enable the model to adapt dynamically to market fluctuations, improving overall alpha generation performance [ong2011].
Diversifying Equities for Growth and Stability
@@ 5 Conclusion
The combined use of the HMM and neural networks in a dual-model approach has shown success in enhancing the accuracy and resilience of trading signals. During the COVID period (2019–2022), the dual-model system achieved an 83% return and a Sharpe ratio of 0.77, highlighting its practical potential.